# On the jump of the cover time in random geometric graphs

### v2

@@ 4. Lower bound on the cover time
Let $\phi(d)=\phi$ be the smallest integer such that any pair of vertices in $G\in\mathcal{G}(n,r,d)$, contained in cubes $\mathrm{Q_{i}}$ and $\mathrm{Q_{j}}$, is not connected by an edge if $|\mathrm{i}-\mathrm{j}|_{1}>\phi$. For example, $\phi(2)=4$. Let $x\in\Lambda_{n}$ be a vertex of $G\in\mathcal{G}(n,r,d)$ with $x\in\mathrm{Q_{j}}$. We say that $x$ is $m$-special if the following two conditions hold:
@@
  • All cubes at distance less or equal to $\phi$ to the pending path in the |.$|.|_{1}$ norm (where the cubes are seen as points of $\mathbb{Z}^{d}$) are empty (see Figure 3 for an example).
@@
In order to make sure that a $\left\lfloor c_{1}\frac{\log n}{r^{d}}\right\rfloor$-pending path completely is contained in $\Lambda_{n}$, we would like its root to be inside $\Lambda_{m_{n}}$, where
@@
Since this holds for any different $x,y\in V^{\prime}$, using Lemma 2.10, by Theorem 2.7, a.a.s. there exist some $\gamma_{1},\gamma_{2}>0$, such that
@@
5. Upper bound on the cover time
@@ 5.1. Random path of good cubes
Assume that $r_{1}=r_{2}$ for the parallelepipeds $\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{T}_{s_{2},r_{2}}$, and assume that $s_{2}>s_{1}$, so that $\mathrm{T}_{s_{2},r_{2}}$ is to the right of $\mathrm{T}_{s_{1},r_{1}}$. We will also throughout this subsection require $s_{2}-s_{1}$ to be odd, so that the parallelepipeds $\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{T}_{s_{2},r_{2}}$ are separated by a an even number of parallelepipeds: we will later lift these restrictions. We now construct a self-avoiding random path $\mathcal{L}_{X}$ (based on the straight path $\mathcal{L}$) of parallelepipeds between $\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{T}_{s_{2},r_{2}}$: let $X$ be a continuous random variable chosen uniformly in $[0,\frac{s_{2}-s_{1}+1}{2}]$, and let $y_{X}\in\Lambda_{n}$ be the point that results from summing $(0,Mr\lceil\alpha\log n\rceil X,0,...,0)$ to the left bottom corner of $\mathrm{T}_{r_{1},\frac{s_{2}+s_{1}+1}{2}}$ (the $x$-coordinate of this point is the midpoint of $\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{T}_{s_{2},r_{2}}$, see Figure  5). Note that we assume here that $y_{X}$ is inside of $\Lambda_{n}$; if this were not the case we subtract $(0,Mr\lceil\alpha\log n\rceil,0,...,0)$ from the midpoint of $\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{T}_{s_{2},r_{2}}$, and the construction that follows is reflected vertically.
@@
We continue in this way until we cross the midpoint (that is, we cross $y_{X}$). Here we continue interpreting the horizontal crossings as paths from left to right but change the orientation of the vertical crossings of good cubes to be from top to bottom. Furthermore, now instead of taking the $\ell$-th vertical crossing of good cubes we take the $(\lceil c\log n\rceil-\ell)$-th vertical crossing of good cubes to create the concatenations (the way of enumerating here is different so that that the concatenations are disjoint for different $\ell$ if we delete the good cubes from $\mathrm{V}^{\prime}(\mathrm{C_{i}})\cup\mathrm{V}^{\prime}(\mathrm{C_{j}})$, see Figure  9). To finish the construction, once we arrive at $\mathrm{T}_{s_{2},r_{2}}$, what is left is the concatenation to the vertical crossing of good cubes $\mathrm{V}^{\prime}(\mathrm{C_{j}})$. Here we proceed just as we did for $\mathrm{C_{i}}$. Select one path that starts from $\mathrm{C_{j}}$ and is made from good cubes of $\mathrm{V}^{\prime}(\mathrm{C_{j}})$ and intersects $\mathcal{Q}_{k-1}$ (the last concatenation available and that passes through $\mathrm{T}_{s_{2},r_{2}}$; there exists at least one such path of good cubes). Concatenate $\mathcal{Q}_{k-1}$ to the last path selected to obtain $\mathcal{Q}_{k}$. Note that the resulting path is self-avoiding by construction, and also note that for different $\ell$, the paths are disjoint if we delete the initial and final good cubes that come from $\mathrm{V}^{\prime}(\mathrm{C_{i}})$ and $\mathrm{V}^{\prime}(\mathrm{C_{j}})$. We denote $\mathcal{Q}_{k}$ by $\mathcal{Q}_{X}^{\ell}$. Delete from $\mathcal{Q}_{k}$ the good cubes that belong to $\mathrm{V}^{\prime}(\mathrm{C_{i}})$ and $\mathrm{V}^{\prime}(\mathrm{C_{j}})$, except for the cubes where the concatenations with $\mathrm{V}^{\prime}(\mathrm{C_{i}})$ and $\mathrm{V}^{\prime}(\mathrm{C_{j}})$ were made at the beginning and the end, and denote the resulting random path by $\mathcal{L}^{\ell}_{X}$ (see Figure  9). In particular, note that the path $\mathcal{L}_{X}^{\ell}$ is a sub-path of $\mathcal{Q}_{X}^{\ell}$ since the latter has $\mathrm{C_{i}}$ and $\mathrm{C_{j}}$ as endpoints (in Figure  9, $\mathcal{Q}_{X}^{\ell}$ corresponds to the blue path together with the green paths in Figure  9 while the path $\mathcal{L}_{X}^{\ell}$ is only the blue path). We point out that the construction of $\mathcal{Q}_{X}^{\ell}$ will be used only in the proof of the upper bound of the cover time above the connectivity threshold in Section  6.
@@ 5.2. Proof of the upper bound on the cover time with fixed radius
Consider then the remaining term. Let $u\in\mathrm{T}_{i,j}$ with $\mathrm{T}_{i,j}\in D_{t}$ and $\{u,v\}\in E$ and $u\notin\mathrm{V}^{\prime}(\mathrm{C_{i}})\cup\mathrm{V}^{\prime}(\mathrm{C_{j}})$. We now claim that the probability that $\mathrm{T}_{i,j}\in D_{t}$, with $1\leq t\leq\frac{s_{2}-s_{1}+1}{2}$, is part of the path of parallelepipeds $\mathcal{L}_{X}$ is bounded above by $\frac{c^{\prime}}{t}$ for some constant $c^{\prime}$ that can be chosen independent of $n$ and $t$ (for $t=0$ we bound this probability by 1). To show this, denote by $y_{X}^{t}$ be the intersection of the segment between the left bottom corner point of $\mathrm{T}_{s_{1},r_{1}}$ and $y_{X}$ and the left line of level $t$ (see Figure  11). By similarity of triangles, $y_{X}^{t}$ is uniform over the left line of level $t$ (see for example the triangles $Ay^{t}_{X}B$ and $Ay_{X}C$ in Figure 11). Now, observe that in order for $\mathrm{T}_{i,j}\in D_{t}$ to be part of $\mathcal{L}_{X}$ the random point $y_{X}^{t}$ has to belong to the left face of $\mathrm{T}_{i,j}$ or to $\mathrm{T}_{i,j-1}$ (the left face of the parallelepiped immediately below, in case such parallelepiped exists): indeed, if $y_{X}^{t}$ were above $\mathrm{T}_{i,j}$, then the parallelepiped does not belong to the path, and if $y^{t}_{X}$ were below $\mathrm{T}_{i,j-1}$, as the angle of the red line is at most $\pi/4$, the red line in Figure 11 does not intersect $\mathrm{T}_{i,j}$. Therefore, the probability that $\mathrm{T}_{i,j}\in D_{t}$ belongs to $\mathcal{L}_{X}$ is bounded from above by the probability that $y_{X}^{t}$ belongs to the left faces of two cubes at level $t$. Since $y_{X}^{t}$ is uniform in its range, then this probability is at most $\frac{2}{t}$. By symmetry, the same is true for $\frac{s_{1}-s_{1}+1}{2}+1\leq t\leq s_{2}-s_{1}$ (for $t=s_{2}-s_{1}+1$ we bound this probability by $1$); in this case we obtain an upper bound of $\frac{2}{s_{2}-s_{1}-t-1}$ for the probability that $\mathrm{T}_{i,j}\in D_{t}$.
@@
We now lift the assumptions on the parallelepipeds: consider first two vertices $x,y$ of $L_{1}$ that are in the same strip of a 2-dimensional slice, but now separated by an odd number of parallelepipeds that are not at the boundary. Let $x\in\mathrm{T}_{s_{1}^{\prime},r_{1}}$ and $y\in\mathrm{T}_{s_{2}^{\prime},r_{1}}$ with $s_{2}^{\prime}-s_{1}^{\prime}$ even such that neither parallelepiped is at the boundary. By the construction given above there exists $x^{\prime}\in L_{1}(G)$ with $x^{\prime}\in\mathrm{T}_{s_{1}^{\prime}+1,r_{1}}$ such that $\mathcal{R}(x^{\prime}\longleftrightarrow y)=O(\log n)$. Note that $d_{E}(x,x^{\prime})=O(\log n)$, since $x$ and $x^{\prime}$ are in adjacent parallelepipeds, and any parallelepiped has logarithmic side length in the first 2 dimensions and fixed side length in the other dimensions. By Lemma  2.12, a.a.s., the graph distance between $x$ and $x^{\prime}$ is then also $O(\log n)$, and thus a.a.s. $\mathcal{R}(x\longleftrightarrow x^{\prime})=O(\log n)$. Then, by application of the triangle inequality (see Lemma 2.4), also a.a.s. $\mathcal{R}(x\longleftrightarrow y)=O(\log n)$. Also, since any two vertices in the same parallelepiped (not necessarily in good cubes, but still in $L_{1}(G)$) satisfy that their graph distance is $O(\log n)$ (see Lemma 2.12, the largest side length of a parallelepiped is $O(\log n)$), by another application of the triangle inequality, also a.a.s. the effective resistance between them is $O(\log n)$. Hence, for any $x^{\prime},y^{\prime}\in L_{1}(G)$ belonging to the largest connected component of the renormalized graph inside the same strip of a $2$-dimensional slice not at the boundary, we have
@@
We now extend this result to larger radii, but still (arbitrarily) smaller than the connectivity threshold:
@@
Let $d\geq 2$, let $G\in\mathcal{G}(n,r,d)$, and let $\varepsilon$ be an arbitrarily small positive constant. Assume furthermore $(1+\varepsilon)r_{g}\leq r\leq(1-\varepsilon)r_{c}$. Then, a.a.s.,
$\tau_{cov}(L_{1}(G))\leq\delta^{\prime}n\log^{2}n,$ (5.11)
where $\delta^{\prime}$ is a constant that only depends on $\varepsilon$.

### v3

@@ 4. Lower bound on the cover time
Let $\phi(d)=\phi$ be the smallest integer such that any pair of vertices in $G\in\mathcal{G}(n,r,d)$, contained in cubes $\mathrm{Q_{i}}$ and $\mathrm{Q_{j}}$, is not connected by an edge if $|\mathrm{i}-\mathrm{j}|_{1}>\phi$. For example, $\phi(2)=3$. Let $x\in\Lambda_{n}$ be a vertex of $G\in\mathcal{G}(n,r,d)$ with $x\in\mathrm{Q_{j}}$. We say that $x$ is $m$-special if the following two conditions hold:
@@
  • All cubes at distance less or equal to $\phi$ to the pending path in the |.$|.|_{1}$ norm (where the cubes are seen as points of $\mathbb{Z}^{d}$) and to the right of $\mathrm{Q_{j}}$ are empty (see Figure 3 for an example).
@@
In order to make sure that a $\left\lfloor c_{1}\frac{\log n}{r^{d}}\right\rfloor$-pending path is completely contained in $\Lambda_{n}$, we would like its root to be inside $\Lambda_{m_{n}}$, where
@@
Since this holds for any different $x,y\in V^{\prime}$, using Lemma 2.10, by Theorem 2.7, a.a.s. there exist some $\delta_{2},\delta_{3}>0$, such that
@@
5. Upper bound on the cover time for fixed radius
@@ 5.1. Random path of good cubes
Assume that $r_{1}=r_{2}$ for the parallelepipeds $\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{T}_{s_{2},r_{2}}$, and assume that $s_{2}>s_{1}$, so that $\mathrm{T}_{s_{2},r_{2}}$ is to the right of $\mathrm{T}_{s_{1},r_{1}}$. We will also throughout this subsection require $s_{2}-s_{1}$ to be odd, so that the parallelepipeds $\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{T}_{s_{2},r_{2}}$ are separated by an even number of parallelepipeds: we will later lift these restrictions. We now construct a self-avoiding random path $\mathcal{L}_{X}$ (based on the straight path $\mathcal{L}$) of parallelepipeds between $\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{T}_{s_{2},r_{2}}$: let $X$ be a continuous random variable chosen uniformly in $[0,\frac{s_{2}-s_{1}+1}{2}]$, and let $y_{X}\in\Lambda_{n}$ be the point that results from summing $(0,Mr\lceil\alpha\log n\rceil X,0,...,0)$ to the left bottom corner of $\mathrm{T}_{r_{1},\frac{s_{2}+s_{1}+1}{2}}$ (the $x$-coordinate of this point is the midpoint of $\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{T}_{s_{2},r_{2}}$, see Figure  5). Note that we assume here that $y_{X}$ is inside of $\Lambda_{n}$; if this were not the case we subtract $(0,Mr\lceil\alpha\log n\rceil,0,...,0)$ from the midpoint of $\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{T}_{s_{2},r_{2}}$, and the construction that follows is reflected vertically.
@@
We continue in this way until we cross the midpoint (that is, we cross $y_{X}$). Here we continue interpreting the horizontal crossings as paths from left to right but change the orientation of the vertical crossings of good cubes to be from top to bottom. Furthermore, now instead of taking the $\ell$-th vertical crossing of good cubes we take the $(\lceil c\log n\rceil-\ell)$-th vertical crossing of good cubes to create the concatenations (the way of enumerating here is different so that that the concatenations are disjoint for different $\ell$ if we delete the good cubes from $\mathrm{V}^{\prime}(\mathrm{C_{i}})\cup\mathrm{V}^{\prime}(\mathrm{C_{j}})$, see Figure  9). To finish the construction, once we arrive at $\mathrm{T}_{s_{2},r_{2}}$, what is left is the concatenation to the vertical crossing of good cubes $\mathrm{V}^{\prime}(\mathrm{C_{j}})$. Here we proceed just as we did for $\mathrm{C_{i}}$. Select one path that starts from $\mathrm{C_{j}}$ and is made from good cubes of $\mathrm{V}^{\prime}(\mathrm{C_{j}})$ and intersects $\mathcal{Q}_{k-1}$ (the last concatenation available and that passes through $\mathrm{T}_{s_{2},r_{2}}$; there exists at least one such path of good cubes). Concatenate $\mathcal{Q}_{k-1}$ to the last path selected to obtain $\mathcal{Q}_{k}$. Note that the resulting path is self-avoiding by construction, and also note that for different $\ell$, the paths are disjoint if we delete the initial and final good cubes that come from $\mathrm{V}^{\prime}(\mathrm{C_{i}})$ and $\mathrm{V}^{\prime}(\mathrm{C_{j}})$. We denote $\mathcal{Q}_{k}$ by $\mathcal{Q}_{X}^{\ell}$. Delete from $\mathcal{Q}_{k}$ the good cubes that belong to $\mathrm{V}^{\prime}(\mathrm{C_{i}})$ and $\mathrm{V}^{\prime}(\mathrm{C_{j}})$, except for the cubes where the concatenations with $\mathrm{V}^{\prime}(\mathrm{C_{i}})$ and $\mathrm{V}^{\prime}(\mathrm{C_{j}})$ were made at the beginning and the end, and denote the resulting random path by $\mathcal{L}^{\ell}_{X}$ (see Figure  9). In particular, note that the path $\mathcal{L}_{X}^{\ell}$ is a sub-path of $\mathcal{Q}_{X}^{\ell}$ since the latter has $\mathrm{C_{i}}$ and $\mathrm{C_{j}}$ as endpoints (in Figure  9, $\mathcal{Q}_{X}^{\ell}$ corresponds to the blue path together with the green paths in Figure  9 while the path $\mathcal{L}_{X}^{\ell}$ is only the blue path). We point out that the construction of $\mathcal{Q}_{X}^{\ell}$ will be used only in the proof of the upper bound of the cover time above the connectivity threshold in Section 6 and Section  7.
@@ 5.2. Proof of the upper bound on the cover time with fixed radius
Consider then the remaining term. Let $u\in\mathrm{T}_{i,j}$ with $\mathrm{T}_{i,j}\in D_{t}$ and $\{u,v\}\in E$ and $u\notin\mathrm{V}^{\prime}(\mathrm{C_{i}})\cup\mathrm{V}^{\prime}(\mathrm{C_{j}})$. We now claim that the probability that $\mathrm{T}_{i,j}\in D_{t}$, with $1\leq t\leq\frac{s_{2}-s_{1}+1}{2}$, is part of the path of parallelepipeds $\mathcal{L}_{X}$ is bounded above by $\frac{c^{\prime}}{t}$ for some constant $c^{\prime}$ that can be chosen independent of $n$ and $t$ (for $t=0$ we bound this probability by 1). To show this, denote by $y_{X}^{t}$ the intersection of the segment between the left bottom corner point of $\mathrm{T}_{s_{1},r_{1}}$ and $y_{X}$ and the left line of level $t$ (see Figure  11). By similarity of triangles, $y_{X}^{t}$ is uniform over the left line of level $t$ (see for example the triangles $Ay^{t}_{X}B$ and $Ay_{X}C$ in Figure 11). Now, observe that in order for $\mathrm{T}_{i,j}\in D_{t}$ to be part of $\mathcal{L}_{X}$ the random point $y_{X}^{t}$ has to belong to the left face of $\mathrm{T}_{i,j}$ or to $\mathrm{T}_{i,j-1}$ (the left face of the parallelepiped immediately below, in case such parallelepiped exists): indeed, if $y_{X}^{t}$ were above $\mathrm{T}_{i,j}$, then the parallelepiped does not belong to the path, and if $y^{t}_{X}$ were below $\mathrm{T}_{i,j-1}$, as the angle of the red line is at most $\pi/4$, the red line in Figure 11 does not intersect $\mathrm{T}_{i,j}$. Therefore, the probability that $\mathrm{T}_{i,j}\in D_{t}$ belongs to $\mathcal{L}_{X}$ is bounded from above by the probability that $y_{X}^{t}$ belongs to the left faces of two cubes at level $t$. Since $y_{X}^{t}$ is uniform in its range, then this probability is at most $\frac{2}{t}$. By symmetry, the same is true for $\frac{s_{1}-s_{1}+1}{2}+1\leq t\leq s_{2}-s_{1}$ (for $t=s_{2}-s_{1}+1$ we bound this probability by $1$); in this case we obtain an upper bound of $\frac{2}{s_{2}-s_{1}-t-1}$ for the probability that $\mathrm{T}_{i,j}\in D_{t}$.
@@
We now lift the assumptions on the parallelepipeds: consider first two vertices $x,y$ of $L_{1}$ that are in the same strip of a 2-dimensional slice, but now separated by an odd number of parallelepipeds that are not at the boundary. Let $x\in\mathrm{T}_{s_{1}^{\prime},r_{1}}$ and $y\in\mathrm{T}_{s_{2}^{\prime},r_{1}}$ with $s_{2}^{\prime}-s_{1}^{\prime}$ even such that neither parallelepiped is at the boundary. By the construction given above there exists $x^{\prime}\in L_{1}(G)$ with $x^{\prime}\in\mathrm{T}_{s_{1}^{\prime}+1,r_{1}}$ such that $\mathcal{R}(x^{\prime}\longleftrightarrow y)=O(\log n)$. Note that $d_{E}(x,x^{\prime})=O(\log n)$, since $x$ and $x^{\prime}$ are in adjacent parallelepipeds, and any parallelepiped has logarithmic side length in the first 2 dimensions and fixed side length in the other dimensions. By Corollary  2.12, a.a.s., the graph distance between $x$ and $x^{\prime}$ is then also $O(\log n)$, and thus a.a.s. $\mathcal{R}(x\longleftrightarrow x^{\prime})=O(\log n)$. Then, by application of the triangle inequality (see Lemma 2.4), also a.a.s. $\mathcal{R}(x\longleftrightarrow y)=O(\log n)$. Also, since any two vertices in the same parallelepiped (not necessarily in good cubes, but still in $L_{1}(G)$) satisfy that their graph distance is $O(\log n)$ (see Corollary 2.12, the largest side length of a parallelepiped is $O(\log n)$), by another application of the triangle inequality, also a.a.s. the effective resistance between them is $O(\log n)$. Hence, for any $x^{\prime},y^{\prime}\in L_{1}(G)$ belonging to the largest connected component of the renormalized graph inside the same strip of a $2$-dimensional slice not at the boundary, we have
@@
6. Upper bound for non-fixed radius
In this section we prove the upperbound for RGGs with radii tending to infinity but below the threshold of connectivity. We begin by noting that the proof given in Section 5 is not enough, as it gives as an upper bound on the resistance of $O(\log n)$, whereas we need an upper bound of magnitude $O(\log(n)/r^{d})$. With this in mind, we therefore introduce a new construction inspired by the one in Section 5 but with refinements that allow for a better flow dissipation. This structure will then also be used in the proof of the upper bound just above the connectivity threshold.
6.1. Resistance bound in the backbone of dense cubes
Recall the definition of $\mu$ introduced in the tessellation of cubes $(\mathrm{Q_{i}})$ in Section 4 of cubes of side length $r/\mu$. In this section, we choose $\mu(n)=\mu+O(1/n)$ to be the smallest real number larger or equal to $\mu$ such that any cube $\mathrm{Q}$ is either completely contained in $\Lambda_{n}$ or it shares at most some faces, so that we avoid boundary issues. With abuse of notation, we denote below $\mu(n)=\mu$. Let $\beta$ be a sufficiently small constant. We say that $\mathrm{Q_{i}}$ is dense if it has at least $\beta r^{d}$ vertices of $G\in\mathcal{G}(n,r,d)$, otherwise we say that it is non-dense.
We see cubes $\mathrm{Q}$ as sites of $\mathbb{Z}^{d}$. By Chernoff’s inequality, the probability that a cube is non-dense is bounded from above by $e^{-\kappa r^{d}}$ for some $\kappa=\kappa(\beta)>0$ (and $\beta$ small enough). We note that $\kappa$ is a fixed constant and does not depend on $r$. Considering dense cubes as occupied sites of a Bernoulli site percolation model, for $r\geq r_{0}$ with $r_{0}$ large enough, we are thus in the super-critical phase of Bernoulli site percolation. That is, a.a.s., there exists a giant component corresponding to dense cubes. We denote by $\mathcal{C}^{site}_{\mathrm{Q}}(n)$ the giant (biggest) cluster of the associated Bernoulli process of dense cubes $\mathrm{Q}$. We prove the following regarding the maximum of the effective resistance in the giant cluster:
@@ 6.1. Resistance bound in the backbone of dense cubes
Let $r\geq r_{0}$ for a sufficiently large constant $r_{0}>0$ and let also $r=O(r_{c})$. Let $x,y\in G\in\mathcal{G}(n,r,d)$ such that $x\in\mathrm{Q_{i}}$ and $y\in\mathrm{Q_{j}}$ and $\mathrm{Q_{i}},\mathrm{Q_{j}}\in\mathcal{C}^{site}_{\mathrm{Q}}(n)$. Then, a.a.s.,
$\mathcal{R}(x\longleftrightarrow y)=O\left(\frac{\log n}{r^{2d}}\right).$